# Liouville polarization and the rigidity of their Lagrangian skeleta in dimension $4$

### v1

@@ Abstract.
In [3], Biran introduced polarizations of closed symplectic manifolds and showed that their Lagrangian skeleta exhibit remarkable rigidity properties. He in particular found that their complements contain only small balls. In this paper, we introduce so-called Liouville polarizations of certain open 4-dimensional symplectic manifolds. This leads to several symplectic embedding results, that in turn lead to new Lagrangian non-removable intersections and a novel phenomenon of Legendrian barriers.
We show for instance that given any connected symplectic 4-manifold $(M,\omega)$ and a 4-ball of smaller volume, their exists an explicit finite union of Lagrangian discs in the 4-ball such that their complement symplectically embeds into $(M,\omega)$, extending a result by Sackel-Song-Varolgunes-Zhu and Brendel.
@@ 1. Introduction and main results
A polarization of a closed symplectic manifold $(M,\omega)$ is a decomposition of the symplectic class $[\omega]$ along positive real multiples of the Poincaré-dual classes of codimension 2 symplectic submanifolds. Closed symplectic manifolds always admit polarizations, and these have proven useful to study the flexible side of symplectic embedding problems [6, 25]. In this work, we introduce polarizations of open symplectic manifolds and show that this concept offers new opportunities, at least in dimension $4$. Referring to §$\S$ 2.3 for the definition, we focus in this introduction on the applications.
@@ 1.1. Symplectic embeddings
A cornerstone of modern symplectic geometry is Gromov’s non-squeezing theorem from [17], stating in dimension four that $B^{4}(a)$ symplectically embeds into $Z^{4}(1)$ only for $a\leq 1$. In [29], Sackel, Song, Varolgunes, and Zhu investigated a refinement of this result, asking how big a set (in the sense of the Hausdorff dimension, for instance) one has to remove from $B^{4}(a)$ such that the complement symplectically embeds into $Z^{4}(1)$. They first proved that it does not suffice to remove a set of lower Minkowski dimension $<2$, and then showed optimality of this result by proving that for $a\leq 2$ it suffices to remove a Lagrangian coordinate disc from $B^{4}(a)$. Later, Brendel in the appendix to [29] showed that for $a<3$ it suffices to remove three Lagrangian pinwheels (certain simple Lagrangian CW-complexes) and a symplectic torus. In their Question 3, these authors then asked:
@@
What is the largest $a$ such that the complement of a 2-dimensional subset of $B^{4}(a)$ symplectically embeds into $Z^{4}(1)$
@@
(ii) A different positive answer to Question 1.1 was independently given by Haim-Kislev, Hind, and Ostrover in [18], and it is interesting to compare the two constructions, see §$\S$ 1.5.
@@
In fact a much more general statement holds. Any affine part of a closed symplectic 4-manifold (defined as the complement of a polarizing divisor, see Definition 2.2) verifies the same kind of flexibility: After removing a suitable Lagrangian CW-complex, it can be embedded into any other affine part of a closed symplectic 4-manifold of larger volume. We shall give the proof of this result in a separate paper, [27], because in addition to the embedding method developped here the proof relies on a relative version of Giroux’s result in [16] and requires specific techniques. In contrast to the results of this paper, the removed CW-complex cannot be explicitly described, however.
@@ 1.2. Rigidity properties of Lagrangian skeleta
Figure 1.1. Examples of grids.
Note that the complement of a grid consists of topological discs. Figure 1.1 shows examples of grids. Note that for two grids $\Gamma$ and $\Gamma^{\prime}$, the Lagrangian CW-complex $\Gamma\times\Gamma^{\prime}$ is smooth if and only if both $\Gamma$ and $\Gamma^{\prime}$ have no vertex, as in (i), (iv), and (vii).
@@ 1.3. Legendrian barriers
Lagrangian CW-complexes  $\Delta$ with symplectically small complement were called Lagrangian barriers in [3] (cf. §$\S$ 1.5). A further consequence of the phenomenon of Lagrangian barriers, that apparently went unnoticed, is a similar Legendrian barrier phenomenon in contact dynamics. A grid is radial if it consists of straight rays emanating from the origin.
@@
(ii) The smooth arcs of $\Lambda_{\delta}$ are indeed Legendrian curves on $S$, because $\iota_{r\partial_{r}}\omega_{{\operatorname{st}}}=\alpha_{{\operatorname{st}}}$ and $\partial_{r}$ is tangent to $\Gamma_{\!\leq\delta_{1}}\!\times\Gamma_{\!\leq\delta_{2}}$. The theorem says that for any Legendrian knot $\Lambda$, every embedded Reeb cylinder of length $\geq\delta_{1}+\delta_{2}$ over $\Lambda$ must intersect $\Lambda_{\delta}$, hence the name Legendrian barrier for $\Lambda_{\delta}$. It follows in particular that for $a>\frac{2}{k}$ there exists no exact contact embedding
@@
In Figure 1.1, all grids are regular except for (iv) and (v) – and possibly (ix).
@@
Theorem 1 is a special case of this theorem, and Theorem 2 is a generalization to a non-compact setting. The exactness in this statement is essential to derive Theorems 4 and 5 on Lagrangian rigidity and Legendrian barriers.
@@ 1.5. Related results and references
Biran’s Lagrangian barriers. As we have already alluded to, the present paper is in many ways a sequel of Biran’s work [3], that is useful to have in mind while reading our paper. We therefore briefly recall it, integrating Giroux’s results from [15, 16] that allow to generalize Biran’s results from the Kähler to the symplectic setting. Let $(M,\omega)$ be a closed symplectic manifold with integral symplectic class (meaning that $[\omega]\in H^{2}(M,\mathbb{Z})$). For every sufficiently large $k$, Donaldson [10] produced a symplectic hypersurface $\Sigma$ Poincaré-dual to $k[\omega]$, called a polarization of degree $k$ of $(M,\omega)$. The complement $M\setminus\Sigma$ was proven to be Weinstein by Giroux [16]. Hence there exists an isotropic CW-complex (called the skeleton) $\Delta\subset M\backslash\Sigma$ to which $M\backslash\Sigma$ retracts. Furthermore, any compact subset of $M\backslash\Delta$ embeds into a ruled symplectic manifold, namely a symplectic sphere bundle over $\Sigma$ whose fibers have area $\frac{1}{k}$, see [2, 3]. A fundamental example to have in mind along this discussion is the following:
@@
([3, §$\S$ 3.1.3]) In the complex projective space $\operatorname{\mathbb{C}P}^{n}$ with Fubini–Study form integrating to 1 over complex lines, the hypersurface
@@
is a symplectic polarization of degree $2k$, and one can arrange the skeleton of $\operatorname{\mathbb{C}P}^{n}\setminus\Sigma_{2k}$ to be the union $\Delta_{2k}(\operatorname{\mathbb{C}P}^{n})$ of $k^{n}$ equally distributed images of $\operatorname{\mathbb{R}P}^{n}$ by projectivized unitary maps. The complement of $2k$ fibers of the symplectic disc bundle over $\Sigma_{2k}$ is symplectomorphic to $B^{2n}(1)\setminus\Delta_{2k}^{n}$, where
@@
Ruled symplectic manifolds are known to have many symplectic invariants bounded by the area of their sphere fibres, which allowed Biran to get interesting non-removable intersections with the skeleton. For instance, for a polarization of degree $k$ with skeleton $\Delta$, every symplectic ball of capacity $\geq\frac{1}{k}$ in $M$ must intersect $\Delta$. He therefore called these skeleta Lagrangian barriers. In particular, such balls cannot be displaced from $\Delta$ by a Hamiltonian isotopy. Combined with the work [9] by Cieliebak and Mohnke, Biran’s decomposition result also implies non-displaceability of small Lagrangian submanifolds from the skeleton. Namely, applying the neck-stretching in [9] to the set-up in [2, 3], one finds that the general Cieliebak–Mohnke Lagrangian width
@@
Let $\Delta$ be the skeleton of a polarization of degree $k$ of $(M,\omega)$. Then for any closed Lagrangian submanifold $L\subset M\backslash\Delta$ it holds that
@@
For instance, for $x\geq\frac{1}{k}$ the product torus $\mathbb{T}^{n}(x)\subset B^{2n}(1)$ whose factors are circles enclosing area $x$ cannot be displaced from $\Delta_{2k}^{n}$ inside $B^{2n}(1)$. Again, Theorem 7 has a Legendrian corollary:
@@
which is the smallest normalized symplectic capacity, and the Cieliebak–Mohnke capacity
@@
is the largest normalized capacity, Theorem 1 states instead that all symplectic capacities of $B^{4}(1)\backslash\Delta_{k}$ are small, and our work [27] generalizes this property to the affine part of any closed symplectic manifold. There are many different capacities, whose estimates have different implications. For instance, applied to the Hofer–Zehnder capacity, Theorem 1 says that the flow of any compactly supported Hamiltonian function on $C^{4}(1)$ that vanishes on a neighbourhood of $\Delta_{k}$ and whose maximum is at least $\frac{2}{k}$ has a non-trivial closed orbit of period $\leq 1$.
Another, more important, gain is that our approach allows to study the rigidity properties of certain singular polarizations. The role of singular polarizations for symplectic embeddings has been observed earlier, and they were used in several works [6, 7, 25, 26]. But the study of their skeleta was left aside in these works. The question of rigidity of Lagrangian skeleta associated to singular polarizations is more subtle than for smooth polarizations. Indeed, the complement of the skeleton of a singular polarization is not ruled anymore, at if its degree is $k$ the Gromov width may not be $\frac{1}{k}$ but may be arbitrarily large [26, Theorem 1]. The present work opens a way to identify singular polarizations for which the complement of the skeleton has small width. It may also be worth noting that some very natural Lagrangian CW-complexes are the skeleta of singular polarizations and not of smooth ones. For instance, the Clifford torus in $\operatorname{\mathbb{C}P}^{2}$ is the skeleton of a polarization by three lines, but cannot be the skeleton of a smooth polarization, see §$\S$ 2.2. And even if the skeleton of a smooth polarization is computable (like $\Delta_{2k}$ for $\Sigma_{2k}$), it may be much easier to obtain it from a singular polarization, see e.g. §$\S$ 3.2.
@@
Theorem 6 versus the main result of [18]. A different positive answer to Question 1.1 was given by Haim-Kislev, Hind, and Ostrover afterwards in [18]. They showed that for every $a$ one can remove from the ball $B^{4}(a)$ a finite number (depending on $a$) of parallel symplectic planes such that the complement symplectically embeds into $Z^{4}(1)$. Their embedding has very different properties from ours: The removed set is symplectic instead of Lagrangian, and the construction is rigid in the sense that one cannot alter the position of the planes, while our grids are rather arbitrary in view of Theorem 6. The key difference is that in contrast to theirs, our embeddings are exact, a property that we crucially need for deriving Theorems 4 and 5.
Organisation of the paper. The paper is organized as follows. In Section 2 we recall the results from [24, 25] on symplectic polarisations in dimension four, introduce Liouville polarizations and explain their main relations with symplectic embeddings. In Section 3 we compute some explicit pairs polarizations/skeleta. Combining these tools, we prove our embedding results in Section 4, deduce the Lagrangian rigidities in Section 5 and the Legendrian ones in Section 6.
@@
[…]
@@
    Usual objects of differential geometry are generalized to multi-curves by just taking a collection of such objects on each component. For instance, a differential $k$-form $\alpha$ on $\Sigma=\cup\Sigma_{i}$ is a collection of differential $k$-forms on the $\Sigma_{i}$. The morphisms between multi-curves $\Sigma,\Sigma^{\prime}$ are smooth maps from $\Sigma$ to $\Sigma^{\prime}$ (in the sense that they are smooth when restricted to each components), and send the regular part and singular locus of $\Sigma$ to those of $\Sigma^{\prime}$. A symplectic embedding between multi-curves is an injective such morphism that pulls back $\omega^{\prime}|_{\Sigma^{\prime}}$ to $\omega|_{\Sigma}$.
@@
    A weighted symplectic multi-curve ${\bf\Sigma}:=\{(\Sigma_{i},\mu_{i})\}$ with normal crossings is a collection of symplectic curves $\Sigma_{i}$ weighted by real numbers $\mu_{i}$, whose total space $\Sigma:=\cup\Sigma_{i}$ is a symplectic multi-curve with normal crossings.
@@
    Given a subset $S$ of a manifold $M$, we write $\operatorname{Op}(S,M)$ or just $\operatorname{Op}(S)$ instead of “some open neighbourhood of $S$ in $M$”.
Acknowledgment. EO thanks Kai Cieliebak for an interesting discussion during a visit to Strasbourg Unversity. FS is grateful to Viktor Ginzburg, Basak Gürel, Marco Mazzucchelli, and Alfonso Sorrentino for organizing the workshop Symplectic Dynamics at Rome in May 2023, where some of the results of this work were presented, and to Umut Varolgunes for many interesting discussions and a great week in June 2023 at Istanbul. FS also cordially thanks Yasha Eliashberg and the Department of Mathematics at Stanford University for their wonderful hospitality during the spring term 2024.
2. Polarizations of symplectic manifolds

### v2

# Liouville polarisations and the rigidity of their Lagrangian skeleta in dimension 44

@@ Abstract.
In 2000, Biran introduced polarisations of closed symplectic manifolds and showed that their Lagrangian skeleta exhibit remarkable rigidity properties. He found, in particular, that their complements contain only small balls. In this paper, we introduce so-called Liouville polarisations of certain open 4-dimensional symplectic manifolds. This leads to several symplectic embedding results, that in turn lead to new Lagrangian non-removable intersections and a novel phenomenon of Legendrian barriers.
We show, for instance, that given any connected symplectic 4-manifold $(M,\omega)$ and a 4-ball of smaller volume, there exists an explicit finite union of Lagrangian discs in the 4-ball such that their complement symplectically embeds into $(M,\omega)$, extending a result by Sackel–Song–Varolgunes–Zhu and Brendel. Other applications are new Lagrangian intersection results and relative versions of the Arnold chord conjecture.
2020 Mathematics Subject Classification. Primary 53D05, Secondary 53D10, 53D12.
Titre. Polarisations de Liouville et rigidité de leurs squelettes lagrangiens en dimension 4
Résumé. En 2000, Biran a introduit les polarisations de variétés symplectiques compactes et a montré que leurs squelettes lagrangiens présentent des propriétés remarquables de rigidité. Il a notamment constaté que leurs complémentaires ne contiennent que des petites boules. Dans cet article, nous introduisons des polarisations dites de Liouville de certaines variétés symplectiques non compactes de dimension 4. Cela conduit à plusieurs résultats de plongements symplectiques, qui à leur tour conduisent à de nouvelles intersections lagrangiennes persistantes et un nouveau phénomène de barrières legendriennes.
Nous montrons par exemple qu’étant donné une variété symplectique connexe $(M,\omega)$ de dimension 4 et une boule de dimension 4 de plus petit volume, il existe une union finie explicite de disques lagrangiens dans la 4-boule dont le complémentaire se plonge symplectiquement dans $(M,\omega)$, ce qui étend un résultat de Sackel–Song–Varolgunes–Zhu et Brendel. D’autres applications sont de nouveaux résultats d’intersections lagrangiennes et des versions relatives de la conjecture des cordes d’Arnold.
@@ 1. Introduction and main results
A polarisation of a closed symplectic manifold $(M,\omega)$ is a decomposition of the symplectic class $\omega]$ along positive real multiples of the Poincaré-dual classes of codimension 2 symplectic submanifolds. Closed symplectic manifolds always admit polarisations, and these have proven useful to study the flexible side of symplectic embedding problems [6, 26]. In this work, we introduce polarisations of open symplectic manifolds and show that this concept offers new opportunities, at least in dimension $4$. Referring to $\S$ 2.3 for the definition, we focus in this introduction on the applications.
@@ 1.1. Symplectic embeddings
A cornerstone of modern symplectic geometry is Gromov’s non-squeezing theorem from [17], stating in dimension four that $B^{4}(a)$ symplectically embeds into $Z^{4}(1)$ only for $a\leq 1$. In [31], Sackel, Song, Varolgunes, and Zhu investigated a refinement of this result, asking how large a set (in the sense of the Hausdorff dimension, for instance) one has to remove from $B^{4}(a)$ such that the complement symplectically embeds into $Z^{4}(1)$. They first proved that it does not suffice to remove a set of lower Minkowski dimension <2<2, and then showed optimality of this result by proving that for $a\leq 2$ it suffices to remove a Lagrangian coordinate disc from $B^{4}(a)$. Later, Brendel in the appendix of [31] showed that for $a<3$ it suffices to remove three Lagrangian pinwheels (certain simple Lagrangian CW-complexes) and a symplectic torus. In their Question 3, these authors then asked:
@@
What is the largest $a$ such that the complement of a 2-dimensional subset of $B^{4}(a)$ symplectically embeds into $Z^{4}(1)$ ?
@@
(ii) A different positive answer to Question 1.1 was independently given by Haim-Kislev, Hind, and Ostrover in [18], and it is interesting to compare the two constructions, see $\S$ 1.5.
@@
This answers in particular the question asked by Claude Viterbo to Paul Biran in 1998, whether the Gromov width of $\mathbb{R}^{4}\setminus(\Gamma\times\Gamma)$ is infinite or not, [34].
@@
In fact a much more general statement holds. Any affine part of a closed symplectic 4-manifold (defined as the complement of a polarizing divisor, see Definition 2.2) verifies the same kind of flexibility: After removing a suitable Lagrangian CW-complex, it can be embedded into any other affine part of a closed symplectic 4-manifold of larger volume. We shall give the proof of this result in a separate paper, [29], because in addition to the embedding method developed here the proof relies on a relative version of Giroux’s result in [16] and requires specific techniques. In contrast to the results of this paper, the removed CW-complex cannot be explicitly described, however.
@@ 1.2. Rigidity properties of Lagrangian skeleta
Figure $1.$ Examples of grids.
Figure 1.1 shows examples of grids. Note that the complement of a grid consists of topological discs, and that for two grids $\Gamma$ and $\Gamma^{\prime}$ the Lagrangian CW-complex $\Gamma\times\Gamma^{\prime}$ is smooth if and only if both $\Gamma$ and $\Gamma^{\prime}$ have no vertex, as in (i) and (iv).
@@ 1.3. Legendrian barriers
Lagrangian CW-complexes  $\Delta$ with symplectically small complement were called Lagrangian barriers in [3] (cf. $\S$ 1.5). A further consequence of the phenomenon of Lagrangian barriers, that apparently went unnoticed, is a similar Legendrian barrier phenomenon in contact dynamics. A grid is radial if it consists of straight rays emanating from the origin.
@@
(ii) The smooth arcs of $\Lambda_{\delta}$ are indeed Legendrian curves on $S$, because $\iota_{r\partial_{r}}\omega_{{\operatorname{st}}}=\alpha_{{\operatorname{st}}}$ and $\partial_{r}$ is tangent to $\Gamma_{\!\leq\delta_{1}}\!\times\Gamma_{\!\leq\delta_{2}}$. The theorem says that for any Legendrian knot $\Lambda$, every embedded Reeb cylinder of length $\geq\delta_{1}+\delta_{2}$ over $\Lambda$ must intersect $\Lambda_{\delta}$, hence the name Legendrian barrier for $\Lambda_{\delta}$. It follows, in particular, that for $a>\frac{2}{k}$ there exists no exact contact embedding
@@
In Figure 1.1, all grids are regular except for (viii) – two of the three non-smooth points of $\Gamma\cup\partial D(A)$ have no Darboux chart as required – and possibly (ix).
@@
Theorem 1 is a special case of this theorem, and Theorem 2 is a generalisation to a non-compact setting. The exactness in this statement is essential to derive Theorems 4 and 5 on Lagrangian rigidity and Legendrian barriers.
@@ 1.5. Related results and references
Biran’s Lagrangian barriers. As we have already alluded to, the present paper is in many ways a sequel of Biran’s work [3], that is useful to have in mind while reading our paper. We therefore briefly recall it, integrating Giroux’s results from [15, 16] that allow to generalise Biran’s results from the Kähler to the symplectic setting. Let $(M,\omega)$ be a closed symplectic manifold with integral symplectic class (meaning that $\omega]\in H^{2}(M,\mathbb{Z})$). For every sufficiently large $k$, Donaldson [10] produced a symplectic hypersurface $\Sigma$ Poincaré-dual to $k[\omega]$, called a polarisation of degree $k$ of $(M,\omega)$. The complement $M\setminus\Sigma$ was proven to be Weinstein by Giroux [16]. Hence there exists an isotropic CW-complex (called the skeleton) $\Delta\subset M\backslash\Sigma$ to which $M\backslash\Sigma$ retracts. Furthermore, any compact subset of $M\backslash\Delta$ embeds into a ruled symplectic manifold, namely a symplectic sphere bundle over $\Sigma$ whose fibers have area $\frac{1}{k}$, see [2, 3]. A fundamental example to have in mind along this discussion is the following:
@@
([3, $\S$ 3.1.3]) In the complex projective space $\operatorname{\mathbb{C}P}^{n}$ with Fubini–Study form integrating to 1 over complex lines, the hypersurface
@@
is a symplectic polarisation of degree $2k$, and one can arrange the skeleton of $\operatorname{\mathbb{C}P}^{n}\setminus\Sigma_{2k}$ to be the union $\Delta_{2k}(\operatorname{\mathbb{C}P}^{n})$ of $k^{n}$ equally distributed images of $\operatorname{\mathbb{R}P}^{n}$ by projectivised unitary maps. The complement of $2k$ fibers of the symplectic disc bundle over $\Sigma_{2k}$ is symplectomorphic to $B^{2n}(1)\setminus\Delta_{2k}^{n}$, where
@@
Ruled symplectic manifolds are known to have many symplectic invariants bounded by the area of their sphere fibres, which allowed Biran to get interesting non-removable intersections with the skeleton. For instance, for a polarisation of degree $k$ with skeleton $\Delta$, every symplectic ball of capacity $\geq\frac{1}{k}$ in $M$ must intersect $\Delta$. He therefore called these skeleta Lagrangian barriers. In particular, such balls cannot be displaced from $\Delta$ by a Hamiltonian isotopy. Combined with the work [9] by Cieliebak and Mohnke, Biran’s decomposition result also implies non-displaceability of small Lagrangian submanifolds from the skeleton. Namely, applying the neck-stretching in [9] to the set-up in [2, 3], one finds that the general Cieliebak–Mohnke Lagrangian width
@@
Let $\Delta$ be the skeleton of a polarisation of degree $k$ of $(M,\omega)$. Then for any closed Lagrangian submanifold $L\subset M\backslash\Delta$ it holds that
@@
See [28, Theorem 3] for a proof. For instance, this theorem combined with a geometric argument explained in [29] implies that for $x\geq\frac{1}{2k}$ and $2k>n$ the product torus $\mathbb{T}^{n}(x)\subset B^{2n}(1)$ whose factors are circles enclosing area $x$ cannot be displaced from $\Delta_{2k}^{n}$ inside $B^{2n}(1)$. Note that for $n=2$ our Theorem 4 gives the same result under the stronger hypothesis $x\geq\frac{1}{k}$.
Again, Theorem 7 has a Legendrian corollary:
@@
which is the smallest normalised symplectic capacity, and the Cieliebak–Mohnke capacity
@@
is the largest normalised capacity, Theorem 1 states instead that all symplectic capacities of $B^{4}(1)\backslash\Delta_{k}$ are small, and our work [29] generalises this property to the affine part of any closed symplectic manifold. There are many different capacities, whose estimates have different implications. For instance, applied to the Hofer–Zehnder capacity, Theorem 1 says that the flow of any compactly supported Hamiltonian function on $C^{4}(1)$ that vanishes on a neighbourhood of $\Delta_{k}$ and whose maximum is at least $\frac{2}{k}$ has a non-trivial closed orbit of period $\leq 1$.
Another, more important, gain is that our approach allows to study the rigidity properties of certain singular polarisations. The role of singular polarisations for symplectic embeddings has been observed earlier, and they were used in several works [6, 7, 26, 27]. But the study of their skeleta was left aside in these works. The question of rigidity of Lagrangian skeleta associated to singular polarisations is more subtle than for smooth polarisations. Indeed, the complement of the skeleton of a singular polarisation is not ruled anymore, and the Gromov width of this complement is not related anymore to the degree but may be arbitrarily close to the Gromov width of the ambient manifold,  [27, Theorem 1]. The present work opens a way to identify singular polarisations for which the complement of the skeleton has small width. It may also be worth noting that some very natural Lagrangian CW-complexes are the skeleta of singular polarisations and not of smooth ones. For instance, the Clifford torus in $\operatorname{\mathbb{C}P}^{2}$ is the skeleton of a polarisation by three lines, but cannot be the skeleton of a smooth polarisation, see $\S$ 2.2. And even if the skeleton of a smooth polarisation is computable (like $\Delta_{2k}$ for $\Sigma_{2k}$), it may be much easier to obtain it from a singular polarisation, see e.g. $\S$ 3.2.
@@
Theorem 6 versus results from [31, 18, 19]. In [31] it is shown that the complement $B^{4}(1)\setminus\Delta_{2}$ of the Lagrangian plane $\Delta_{2}$ symplectically embeds into $Z^{4}(\frac{1}{2})$. While Theorems 1 and 6 do not provide such an embedding, a modification of our construction yields a symplectic embedding $B^{4}(1)\setminus\Delta_{2}\to Z^{4}(\frac{1}{2}+\varepsilon)$ for every $\varepsilon>0$, see Remark 4.9.
A different positive answer to Question 1.1 was given by Haim-Kislev, Hind, and Ostrover afterwards in [18]. They showed that for every $a$ one can remove from the ball $B^{4}(a)$ a finite number (depending on $a$) of parallel symplectic planes such that the complement symplectically embeds into $Z^{4}(1)$. Their embedding has very different properties from ours: The removed set is symplectic instead of Lagrangian, and the construction is rigid in the sense that one cannot alter the position of the planes, while our grids are rather arbitrary in view of Theorem 6. The key difference is that in contrast to theirs, our embeddings are exact, a property that we crucially need for deriving Theorems 4 and 5.
In [19], the same authors compute the cylindrical capacity of the complement $B^{4}(1)\setminus P_{t}$ of a plane with Kähler angle $t\in[0,1]$ to be $\frac{1+t}{2}$, interpolating between the previously known case of the Lagrangian plane $\Delta_{2}=P_{0}$ from [31] and the flexible case of a complex plane. By contrast, our embedding method leads only to barriers which are Lagrangian skeleta. It would be interesting to know whether these Lagrangian skeleta also belong to natural 1-parameter families of sets which do not remain Lagrangian, but still retain some barrier property.
Organisation of the paper. The paper is organised as follows. In Section 2 we recall the results from [25, 26] on symplectic polarisations in dimension four, introduce Liouville polarisations and explain their main relations with symplectic embeddings. In Section 3 we compute some explicit pairs polarisations/skeleta. Combining these tools, we prove our embedding results in Section 4, deduce the Lagrangian rigidities in Section 5 and the Legendrian ones in Section 6.
@@
    A symplectic embedding $\varphi\colon(M,d\alpha)\to(M^{\prime},d\alpha^{\prime})$ between exact symplectic manifolds is $(\alpha,\alpha^{\prime})$-exact if $\varphi^{*}\alpha^{\prime}=\alpha+df$ for a smooth function $f$ on $M$. Equivalently, $\int_{\gamma}\alpha=\int_{\varphi(\gamma)}\alpha^{\prime}$ for all closed curves $\gamma$ in $M$. It is enough to check this equality of actions on a set of closed curves that generate $H_{1}(M;\mathbb{R})$. If $H^{1}(M;\mathbb{R})=0$, then $(\alpha,\alpha^{\prime}$)-exactness is automatic; and otherwise, the notion of $(\alpha,\alpha^{\prime}$)-exactness may depend on the choice of primitives $\alpha,\alpha^{\prime}$. We sometimes write $\lhook\joinrel\xrightarrow{\;\!\alpha_{{\operatorname{st}}}\,}$ to abbreviate “there exists an $(\alpha_{{\operatorname{st}}},\alpha_{{\operatorname{st}}})$-exact symplectic embedding”.
@@
    Usual objects of differential geometry are generalised to multi-curves by just taking a collection of such objects on each component. For instance, a differential $k$-form $\alpha$ on $\Sigma=\bigcup\Sigma_{i}$ is a collection of differential $k$-forms on the $\Sigma_{i}$. The morphisms between multi-curves $\Sigma,\Sigma^{\prime}$ are smooth maps from $\Sigma$ to $\Sigma^{\prime}$ (in the sense that they are smooth when restricted to each components), and send the regular part and singular locus of $\Sigma$ to those of $\Sigma^{\prime}$. A symplectic embedding between multi-curves is an injective such morphism that pulls back $\omega^{\prime}|_{\Sigma^{\prime}}$ to $\omega|_{\Sigma}$.
@@
    A weighted symplectic multi-curve ${\bf\Sigma}:=\{(\Sigma_{i},\mu_{i})\}$ with normal crossings is a collection of symplectic curves $\Sigma_{i}$ weighted by real numbers $\mu_{i}$, whose total space $\Sigma:=\bigcup\Sigma_{i}$ is a symplectic multi-curve with normal crossings.
@@
    Given a subset $S$ of a manifold $M$, we write $\operatorname{Op}(S,M)$ or just $\operatorname{Op}(S)$ instead of “some open neighbourhood of $S$ in MM”.
Acknowledgment. EO thanks Kai Cieliebak for an interesting discussion during a visit to Strasbourg University. FS is grateful to Viktor Ginzburg, Basak Gürel, Marco Mazzucchelli, and Alfonso Sorrentino for organizing the workshop Symplectic Dynamics at Rome in May 2023, where some of the results of this work were presented, and to Umut Varolgunes for many interesting discussions and a great week in June 2023 at Istanbul. We also thank Umut for pointing out a gap in an earlier version, and Paul Biran for pointing out reference [34]. FS also cordially thanks Yasha Eliashberg and the Department of Mathematics at Stanford University for their wonderful hospitality during the spring term 2024.
2. Polarisations of symplectic manifolds